x-3y=6
3x+y=-22
its using cramer's rule
i've tried to work the problem but i jus keep getting 46 and 40

Answer by nerdybill(5413)

x-3y=6
3x+y=-22
.
coefficient matrix:
1 -3
3 1
Det of coefficients: (1-(-9)) = 10
.
x-matrix
6 -3
-22 1
Detx: 6 - 66 = -60
.
y-matrix
1 6
3 -22
Dety: -22 - 18 = -40
.
x = Detx/Detcoef = -60/10 = -6
y = Dety/Detcoef = -40/10 = -4
.